# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### 28 Feb 2026 v4

@@ 3.1 Neighborhood bias
Fix $j\in[n]$. Let $M^{(-j)}_{\mathrm{embed}}$ and $M^{(-j)\star}_{\mathrm{embed}}$ be the $n\times n$ embeddings of $M^{(-j)}$ and $M^{(-j)\star}$ defined as in Lemma 3.1. Then
By Lemma 3.1,
$\sup_{1\leq j\leq n}\|M^{(-j)}_{\mathrm{embed}}-M\|_{\max}\leq\frac{2}{n}.$
Next, since
$M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)\star}_{\mathrm{embed}}=\mathbb{E}[M^{(-j)}_{\mathrm{embed}}\mid P],$
we have
$M^{\star}-M^{(-j)\star}_{\mathrm{embed}}=\mathbb{E}[M-M^{(-j)}_{\mathrm{embed}}\mid P],$
and thus, by Jensen’s inequality and Lemma 3.1,
$\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\mathbb{E}\big[\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\mid P\big]\leq\frac{2}{n}.$
Combining the bounds,
$\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\|M-M^{\star}\|_{\max}+\frac{4}{n}.$
Taking the supremum over $j$ yields (3.4). ∎
Proposition 3.3 (Theorem 1 of Zhang et al. (2017): Row-wise risk).
Work under the assumptions of Zhang et al. (2017) and choose the ZLZ quantile cutoff parameter $h_{\mathrm{ZLZ}}=C_{0}(\log n/n)^{1/2}$, so that the resulting neighborhood sizes satisfy $|N_{i}|\geq(n\log n)^{1/2}$ (Lemma 2(1) in their supplementary material). Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the two-hop matrix $M=A^{2}/n$ via their distance-and-quantile rule, and define
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$
$\frac{1}{n}\sum_{j=1}^{n}\big(\widetilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$
Proof.
This is exactly the row-wise high-probability bound proved in the supplementary material of Zhang et al. (2017): after the decomposition (1), the row-averages of the two terms are bounded in (15) and (16) of their supplementary material, uniformly for all $i$ on an event of probability $1-O(n^{-c})$. ∎
Fix $j\in[n]$. Let $A^{(-j)}$ be the adjacency matrix obtained by deleting row and column $j$ from $A$, and let $\widetilde{P}$ denote the ZLZ neighborhood-smoothing estimator computed from $A^{(-j)}$, with neighborhoods $N_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $(A^{(-j)})^{2}/(n-1)$. Recall (3.1). For each $i\neq j$ and $\ell\neq j$, define
$U^{(-j)}_{i\ell}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(A_{k\ell}-P_{k\ell}),\qquad B^{(-j)}_{i\ell}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(P_{k\ell}-P_{i\ell}),$
so that
$\widetilde{P}_{i\ell}-P_{i\ell}=U^{(-j)}_{i\ell}+B^{(-j)}_{i\ell}$.
Then for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2}.$
Moreover, applying Theorem 1 of Zhang et al. (2017) to the (n1)(n-1)-node graph $A^{(-j)}$ yields, with probability at least $1-c_{1}n^{-c_{2}}$, uniformly for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Identifying ZLZ’s notation via
$J_{1}^{(-j)}(i,\ell):=\big(U^{(-j)}_{i\ell}\big)^{2},\qquad J_{2}^{(-j)}(i,\ell):=\big(B^{(-j)}_{i\ell}\big)^{2},$
the ZLZ bounds (with (n1)(n-1) in place of $n$) give
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Proof.
The decomposition
$\widetilde{P}_{i\ell}-P_{i\ell}=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\bigl[(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\bigr]$
is identical to ZLZ equation (1), and the inequality
$(\widetilde{P}_{i\ell}-P_{i\ell})^{2}\leq 2J_{1}^{(-j)}(i,\ell)+2J_{2}^{(-j)}(i,\ell)$
follows exactly as in their argument, with the identifications $J_{1}^{(-j)}(i,\ell)=(U^{(-j)}_{i\ell})^{2}$ and $J_{2}^{(-j)}(i,\ell)=(B^{(-j)}_{i\ell})^{2}$. Next, ZLZ’s control of two-hop quantities
$\frac{1}{n}(A^{2})_{ab}-\frac{1}{n}(P^{2})_{ab}$
is established in equations (4)-(6) of their supplementary material. Lemma 3.1 shows that the same bounds hold when $A$ is replaced by $A^{(-j)}$ (and $n$ by n1n-1). Therefore, all subsequent ZLZ steps that rely on these concentration inputs remain valid for the leave-one-out graph.
ZLZ bound the row-average of $J_{1}$ using their decomposition (11) and bounds (12), (14), and (15). The same argument applies verbatim to $A^{(-j)}$ and $N_{i}^{(-j)}$, yielding
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Similarly, ZLZ bound the row-average of $J_{2}$ using Cauchy-Schwarz and their “claim 2 of Lemma 2,” concluding in equation (16). Again, Lemma 3.1 ensures the same concentration inputs hold for $A^{(-j)}$, giving
$\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Combining these bounds yields the stated row-wise risk bound for the leave-one-out estimator. ∎
Assumption 3.5 (Piecewise bi-Lipschitz smoothness).
The graphon $W:[0,1]^{2}\to[0,1]$ admits a finite vertical partition $0=b_{0}<b_{1}<\cdots<b_{K}=1$ with blocks $I_{m}:=(b_{m},b_{m+1})$ such that:
  1. 1.
    $|W(u,x)-W(u^{\prime},x)|\leq L|u-u^{\prime}|.$
  2. 2.
    $\lambda\big(\{x\in[0,1]:|W(u,x)-W(u^{\prime},x)|\geq\eta\}\big)\geq\delta.$
  3. 3.
    $\int_{0}^{1}(W(u,x)-W(u^{\prime},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|u-u^{\prime}|^{2}.$
Remark 3.6 (Scope of Assumption 3.5.).
Remark 3.7 (Good-node and good-edge sets.).
Fix $\varepsilon>0$ satisfying the feasibility condition
$b_{m}+\varepsilon<b_{m+1}-\varepsilon\qquad\text{for all }m=0,\dots,K-1,$
so that each interior interval $b_{m}+\varepsilon,b_{m+1}-\varepsilon]$ is nonempty. Define the block interior
$I_{\varepsilon}:=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,\;b_{m+1}-\varepsilon].$
Let the random good-node set be
$\mathcal{G}_{\varepsilon}:=\{i\in[n]:\ \xi_{i}\in I_{\varepsilon}\},$
and the good-edge set
$\mathcal{E}_{\varepsilon}:=\{(i,j):i\neq j,\ i\in\mathcal{G}_{\varepsilon},\ j\in\mathcal{G}_{\varepsilon}\}.$
Since $\xi_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$, we have $|\mathcal{G}_{\varepsilon}|\sim\mathrm{Bin}\big(n,\lambda(I_{\varepsilon})\big)$. In particular, for any $t>0$ Hoeffding’s inequality gives
$\mathbb{P}\!\left(\left|\frac{|\mathcal{G}_{\varepsilon}|}{n}-\lambda(I_{\varepsilon})\right|\geq t\right)\leq 2\exp(-2nt^{2}).$
Consequently, for fixed $\varepsilon>0$ the fraction of good nodes satisfies $|\mathcal{G}_{\varepsilon}|/n=\lambda(I_{\varepsilon})+O_{\mathbb{P}}(n^{-1/2})$, and if $\lambda(I_{\varepsilon})=1-O(\varepsilon)$ then $|\mathcal{G}_{\varepsilon}|=n(1-O(\varepsilon))$ with exponentially high probability.
Lemma 3.8 (Rowwise $L^{2}$ closeness implies latent closeness).
Assume Assumption 3.5. Fix $\varepsilon>0$ and define
$\mathcal{I}_{\varepsilon}=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,b_{m+1}-\varepsilon].$
Let $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$ and set $P_{ab}=W(U_{a},U_{b})$. Let $r_{n}=(\log n/n)^{1/4}$. There exist constants $c_{0},c_{1}>0$ depending only on Assumption 3.5 and $\varepsilon$ such that the following holds. If
$M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq c_{0},\qquad\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},$
then with probability at least $1-n^{-c}$ (for some $c>0$) and for all sufficiently large $n$,
$U_{i},U_{k}\text{ lie in the same block},\qquad|U_{i}-U_{k}|\leq c_{1}\sqrt{M_{n}}\,r_{n}.$
Proof.
Let $f_{i}(x)=W(U_{i},x)$ and $f_{k}(x)=W(U_{k},x)$, and define
$D(U_{i},U_{k})=\int_{0}^{1}(f_{i}(x)-f_{k}(x))^{2}\,dx.$
Concentration of the empirical row distance. Conditional on $(U_{i},U_{k})$, the variables
$Z_{\ell}:=(P_{i\ell}-P_{k\ell})^{2}=(W(U_{i},U_{\ell})-W(U_{k},U_{\ell}))^{2}$
are i.i.d. in $\ell$, bounded in $[0,1]$, and satisfy $\mathbb{E}[Z_{\ell}\mid U_{i},U_{k}]=D(U_{i},U_{k})$. Hoeffding’s inequality gives, for any $t>0$,
$\mathbb{P}\!\left(\left|\frac{1}{n}\sum_{\ell=1}^{n}Z_{\ell}-D(U_{i},U_{k})\right|\geq t\ \middle|\ U_{i},U_{k}\right)\leq 2e^{-2nt^{2}}.$
Taking $t=C(\log n/n)^{1/2}$ yields that, with probability at least $1-n^{-c}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq D(U_{i},U_{k})-C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.5)
If $U_{i},U_{k}$ lie in different blocks, Assumption 3.5(ii) gives $D(U_{i},U_{k})\geq\eta^{2}\delta$. Using (3.5), for large $n$,
Choosing $c_{0}=\eta^{2}\delta/4$ makes this incompatible with the assumed upper bound. Thus $U_{i},U_{k}$ must lie in the same block.
On the event of $U_{i},U_{k}$ lie in the same block and belong to $\mathcal{I}_{\varepsilon}$, so Assumption 3.5(iii) implies
$D(U_{i},U_{k})\geq\kappa(\varepsilon)|U_{i}-U_{k}|^{2}.$
Combining this with (3.5) and the assumed upper bound gives
$\kappa(\varepsilon)|U_{i}-U_{k}|^{2}\leq\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}+C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq(M_{n}+C)\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Therefore,
$|U_{i}-U_{k}|\leq\sqrt{\frac{M_{n}+C}{\kappa(\varepsilon)}}\,\Bigl(\frac{\log n}{n}\Bigr)^{1/4}\leq c_{1}\sqrt{M_{n}}\,r_{n},$
after adjusting constants. ∎
Corollary 3.9 (Bias bound on the good set).
Fix $\varepsilon>0$ satisfying the feasibility condition above and suppose the neighborhood regularity condition (display (3.4) in the main text) holds with constant $C_{2}$. Then there exists a constant $C(\varepsilon,C_{2})>0$ (depending on $L,\eta,\delta,\kappa(\varepsilon)$ and $C_{2}$) such that, with probability at least $1-c_{1}n^{-c_{2}}$,
$\sup_{\begin{subarray}{c}i\neq j:\\ (i,j)\in\mathcal{E}_{\varepsilon}\end{subarray}}|B_{ij}^{(-j)}|=O_{\mathbb{P}}\big(C(\varepsilon,C_{2})\,r_{n}\big),\qquad r_{n}=(\log n/n)^{1/4}.$
Assumption 3.10 (Neighborhood regularity).
For each $(i,j)\in\mathcal{E}_{\varepsilon}$, the leave-one-out neighborhood $\mathcal{N}_{i}^{(-j)}$ satisfies
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}$
with probability at least $1-n^{-c}$.
Remark.
Assumption 3.10 is consistent with the rowwise risk bounds of Zhang-Levina-Zhu together with the two-hop stability lemmas established above. We do not provide a fully explicit derivation of this condition for LOO neighborhoods here, but it is compatible with existing theory.
Lemma 3.11 (Bias bound for the LOO neighborhood smoother).
Assume Assumption 3.5 and let $r_{n}=(\log n/n)^{1/4}$. Suppose that, with probability at least $1-n^{-c}$, for all $i\neq j$,
$\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq C_{2}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.6)
Then for all $i\neq j$ with $U_{i}\in\mathcal{I}_{\varepsilon}$,
$|B^{(-j)}_{ij}|\leq Cr_{n}.$
The canonical choice $h_{n}\asymp\sqrt{n\log n}$ balances the stochastic fluctuation scale $h_{n}^{-1/2}$ and the bias scale $r_{n}=(\log n/n)^{1/4}$. Undersmoothing ($h_{n}=o(\sqrt{n\log n})$) reduces bias at the cost of larger variance, while oversmoothing increases bias. Thus the canonical scaling is optimal for first-order mean-squared error and yields first-order coverage up to the explicit bias cushion.
Proof.
Fix $i\neq j$ and write $f_{u}(x)=W(u,x)$. Let $N=N_{i}^{(-j)}$ and define
$D(u,u^{\prime})=\int_{0}^{1}(f_{u}-f_{u^{\prime}})^{2}.$
First we show that conditional on $(U_{i},U_{k})$, the variables $(P_{i\ell}-P_{k\ell})^{2}$ are i.i.d. in $\ell$, bounded in $[0,1]$, and have mean $D(U_{i},U_{k})$. Hoeffding’s inequality and a union bound over $k\in N$ (using $|N|\leq h_{n}\asymp\sqrt{n\log n}$) give that with probability $1-n^{-c}$,
$\sup_{k\in N}\left|\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}-D(U_{i},U_{k})\right|\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.7)
Combining (3.6) and (3.7),
$\frac{1}{|N|}\sum_{k\in N}D(U_{i},U_{k})\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.8)
Now we control the average latent distance between $U_{i}$ and its neighbors in $N$. Let
$\mathcal{C}_{i}=\{k\in N:U_{k}\text{ lies in a different block than }U_{i}\}.$
Assumption 3.5(ii) gives $D(U_{i},U_{k})\geq\eta^{2}\delta$ for $k\in\mathcal{C}_{i}$. Using (3.8),
Decompose
$\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}=\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}+\frac{1}{|N|}\sum_{k\in\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}.$
The second term is at most $|\mathcal{C}_{i}|/|N|$. For $k\in N\setminus\mathcal{C}_{i}$ (same block), Assumption 3.5(iii) gives $D(U_{i},U_{k})\geq\kappa(\varepsilon)|U_{i}-U_{k}|^{2}$, hence
$\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}\leq\frac{1}{\kappa(\varepsilon)}\cdot\frac{1}{|N|}\sum_{k\in N}D(U_{i},U_{k}).$
Using (3.8) and $(\dagger)$,
$\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.8)
Step 3: Bias bound. Lipschitz continuity in the first argument gives
$|P_{kj}-P_{ij}|\leq L|U_{k}-U_{i}|.$
Thus
$|B^{(-j)}_{ij}|=\left|\frac{1}{|N|}\sum_{k\in N}(P_{kj}-P_{ij})\right|\leq L\sqrt{\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}}\leq Cr_{n},$
using (3.8) and $r_{n}=(\log n/n)^{1/4}$. ∎
Lemma 3.11 establishes a bias bound for the leave-one-out neighborhood smoother, showing that under the piecewise bi-Lipschitz smoothness assumption, the bias term $\sup_{\begin{subarray}{c}i\neq j:\\ (i,j)\in\mathcal{E}_{\varepsilon}\end{subarray}}|B_{ij}^{(-j)}|$ is of order $r_{n}=(\log n/n)^{1/4}$ for nodes $i$ and $j$ whose latent positions lie in the interior of the same block. This result relies on the control of the average two-hop distance between node $i$ and its neighbors in the leave-one-out neighborhood, as well as on the Lipschitz continuity of the graphon within blocks.
Remark 3.12.
Under the canonical choice $h_{n}\asymp\sqrt{n\log n}$, the stochastic fluctuation scale $h_{n}^{-1/2}$ and the bias scale $r_{n}=(\log n/n)^{1/4}$ are of the same order. Thus the normal approximation is first-order accurate, and bias cannot be ignored without undersmoothing. Choosing a smaller bandwidth would reduce bias at the cost of increased variance, while larger bandwidths inflate bias. The canonical scaling balances these two effects.
3.2 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
We introduce the conditional variance for $U_{ij}^{(-j)}$:
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.9)
so that $V_{ij}\leq 1/(4h)$. While $V_{ij}$ is analytically useful and measurable with respect to the $\sigma$-field generated by $(\mathcal{N}_{i}^{(-j)},P)$, it depends on the unknown latent probabilities $\{P_{kj}\}$.
To obtain tight, variance-adaptive control of $U_{ij}^{(-j)}$ using only observable quantities, we define the (Bessel-corrected) sample variance of the edges in the leave-one-out neighborhood:
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}),$ (3.10)
where $\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical mean.
Instead of relying on classical Bernstein bounds that require the true variance $V_{ij}$, we utilize the Empirical Bernstein inequality (Theorem 11 of Maurer and Pontil (2009)), which directly incorporates the sample variance $\hat{\sigma}_{ij}^{2}$.
For the normal-approximation intervals we estimate the conditional variance
$V_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})$
by the plug-in estimator
$\widehat{V}_{ij}:=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}),$
which replaces the unknown probabilities $P_{kj}$ by their smoothed estimates.
Lemma 3.13 (Empirical Bernstein inequality).
Assume $h_{n}\geq 2$. For any $\delta\in(0,1)$,
$\mathbb{P}\!\left(|U_{ij}^{(-j)}|\geq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\ln(2/\delta)}{h_{n}}}+\frac{7\ln(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\leq\delta.$
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Moreover, by the leave-one-out construction, $\mathcal{N}_{i}^{(-j)}$ depends only on $A^{(-j)}$ and is independent of $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ given $P$.
Their empirical mean is
$\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},$
with conditional expectation
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}.$
Thus $U_{ij}^{(-j)}$ is exactly the deviation of the empirical mean from its conditional expectation.
Applying Theorem 11 of Maurer and Pontil (2009) to the variables $X_{k}:=A_{kj}$ and again to $1-X_{k}$ yields the stated two-sided inequality. ∎
Lemma 3.13 provides a strictly finite-sample, variance-adaptive control of $U_{ij}^{(-j)}$ directly in terms of the empirical variance $\hat{\sigma}_{ij}^{2}$, thereby bypassing the need for asymptotic plug-in arguments.
3.3 Berry-Esseen normal approximation
We now establish a normal approximation for the variance component $U_{ij}^{(-j)}$ defined in (3.2). Recall
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$
and define the conditional variance
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}\mid\mathcal{N}_{i}^{(-j)},P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
so that
$V_{ij}=\mathrm{Var}(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P)=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2}.$
Under Assumption 2.2, there exists $\eta\in(0,1/2)$ such that

### 14 Mar 2026 v5

@@ 3.1 Finite-Sample Confidence Intervals via Empirical Bernstein Bounds
Let $E_{1}$ be the event that the empirical Bernstein bound holds, which has probability $\mathbb{P}(E_{1})\geq 1-\alpha$ by Lemma 3.1. Let $E_{2}$ be the event that the uniform bias bound holds over $\mathcal{E}_{\varepsilon}$, which has probability $\mathbb{P}(E_{2})\geq 1-n^{-c}$ by Corollary 4.13. On the intersection $E_{1}\cap E_{2}$, the distance from the estimator $\tilde{P}_{ij}$ to the neighborhood expected value is bounded by $w_{ij}^{(B)}(\alpha)$, and the distance from the neighborhood expected value to the true $P_{ij}$ is bounded by $Cr_{n}$. Applying the triangle inequality via the Minkowski sum, $P_{ij}$ must lie in $\mathrm{CI}^{(B)}_{ij}(\alpha)\oplus[-Cr_{n},\;Cr_{n}]$. By the union bound, $\mathbb{P}(E_{1}\cap E_{2})=1-\mathbb{P}(E_{1}^{c}\cup E_{2}^{c})\geq 1-\alpha-n^{-c}$. ∎
Remark 3.4 (Advantage of the Empirical Bound).
Unlike standard concentration inequalities that require substituting an estimated variance $\widehat{V}_{ij}$ for the true variance $V_{ij}$-a step that forces reliance on Slutsky’s Theorem and asymptotic normality-the Empirical Bernstein interval completely avoids the cost of studentization. The interval $\mathrm{CI}_{ij}^{(B)}(\alpha)$ automatically adapts to the empirical variance of the neighborhood while strictly preserving its $1-\alpha$ coverage probability for the localized average at any finite sample size.
3.2 Asymptotic Normality and Normal-Approximation Intervals
While the Empirical Bernstein bounds provide strictly finite-sample guarantees, it is also of fundamental theoretical and practical interest to establish the asymptotic distribution of the localized average. Under the non-degeneracy condition (Assumption 2.2), the conditional variance of the LOO summands is bounded uniformly away from zero. This allows us to establish a uniform Berry-Esseen bound for the standardized stochastic fluctuation.
Recall that the stochastic fluctuation term is defined as the average of the mean-zero noise over the leave-one-out neighborhood:
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}).$ (3.9)
Conditional on the neighborhood $\mathcal{N}_{i}^{(-j)}$ and the latent probability matrix $P$, the edge indicators $A_{kj}$ are independent Bernoulli random variables. Therefore, the conditional variance of $U_{ij}^{(-j)}$ is strictly diagonal and is given by
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}).$ (3.10)
To facilitate the subsequent Berry-Esseen analysis, we define the unscaled variance of the neighborhood sum, $\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$, as
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.11)
so that $V_{ij}=\sigma_{ij}^{2}/h_{n}^{2}$. Under Assumption 2.2, there exists $\eta\in(0,1/2)$ such that